# ORACLE: Explaining Feature Interactions in Neural Networks with ANOVA

### v2

@@ A.5 Interaction strength consistency and Top-$K$ / NDCG@K alignment
We finally connect Top-$K$ recovery to the ranking metric NDCG@K. Let $G^{\star}_{jk}$ denote the oracle “gain” associated with pair $(j,k)$, for example $G^{\star}_{jk}=S^{\star}_{jk}$ or a monotone transform thereof. Given a permutation $\pi$ of the pairs, the (unnormalized) DCG@K of $\pi$ with respect to $G^{\star}$ is
@@ A.7 Further theoretical remarks: grid resolution, misspecification, and higher-order terms
All oracle quantities in Appendix A.2 are defined as orthogonal projections in the Hilbert space $L^{2}(P_{X})$ of the trained predictor $f$. In particular, the main-effect and interaction components $(m_{j},g_{jk})$ and the residual $r$ are defined by their $L^{2}(P_{X})$-best approximation properties, regardless of whether $f$ arises from a “correctly specified” probabilistic model.
@@
  • When the true mechanism contains higher-order or non-factorizable structure, these contributions are distributed across the pairwise surfaces $g_{jk}$ and the residual $r$ in the way that minimizes $L^{2}(P_{X})$ error. Our oracle is therefore “model-centric”: it asks how $f$ can be best approximated by main and pairwise terms under the observed $P_{X}$, rather than asserting that $g_{jk}$ coincides with a causal or structural notion of two-way interaction.
@@ Appendix B Additional implementation details
After fitting the factorial surrogate in the ANOVA basis, we apply a lightweight post-processing step to enforce discrete ANOVA-style constraints on the main-effect and interaction tables. This step consists of (i) centering each main-effect table to have zero mean over its levels, (ii) centering each pairwise interaction table so that all row and column sums vanish, and (iii) absorbing any constant offsets induced by these operations into the global mean $\hat{\mu}$. The procedure is summarized in Algorithm 1; it requires only a few passes over the discrete tables and is negligible compared to training the backbone or fitting the surrogate.
Algorithm 1 Residual centering and $\mu$-rebalancing of ANOVA tables
@@
1:// Center main effects
@@
11:  $\bar{g}_{jk}\leftarrow\frac{1}{|\mathcal{B}_{j}||\mathcal{B}_{k}|}\sum_{\ell_{j},\ell_{k}}\hat{g}_{jk}(\ell_{j},\ell_{k})$
@@
14:  subtract row means so that $\sum_{\ell_{j}}\hat{g}_{jk}(\ell_{j},\ell_{k})=0$ for all $\ell_{k}$
15:  subtract column means so that $\sum_{\ell_{k}}\hat{g}_{jk}(\ell_{j},\ell_{k})=0$ for all $\ell_{j}$
@@
17: Centered tables $\hat{\mu}$, $\{\hat{m}_{j}\}$, $\{\hat{g}_{jk}\}$ satisfying discrete ANOVA-style constraints
@@ Appendix D Classical Main-Effect and Interaction Plots
ANOVA.
Table 9 reports an ANOVA-style variance decomposition of the ORACLE surrogate trained on Airfoil (Backbone A). Among the main effects, Frequency and Suction side dominate the contribution to model variation: their sums of squares are roughly $1.4\times 10^{4}$ and $8.5\times 10^{3}$, respectively, clearly larger than those of Chord length, Angle of attack, and Free stream. In the normalized %Total column, these two factors alone account for about $18.2\%$ and $11.1\%$ of the overall ORACLE importance, while the remaining main effects are comparatively modest. On the interaction side, the pairs Frequency:Chord length, Frequency:Suction side, Angle of attack:Suction side, and Angle of attack:Chord length exhibit particularly large sums of squares and %Total values in the mid single digits, revealing strong non-additive structure between these factors. Interactions involving Free stream (e.g., Frequency:Free stream, Chord length:Free stream) are present but comparatively weaker. The residual term still carries about $26.3\%$ of the normalized total contribution, which indicates that a pure two-way ORACLE decomposition does not capture all higher-order or fine-grained nonlinear effects, and that the table is best viewed as a screening device for the most influential factors and pairs.
From a practical perspective, we recommend using the ANOVA-style table in Table 9 as a first-stage diagnostic before examining the classical plots shown later in this section.
  1. 1.
    Ranking rows by Sum Sq or the normalized %Total quickly identifies the main effects and interactions that account for most of the ORACLE importance (here, Frequency, Suction side, and their interactions).
  2. 2.
    The %Expl. column rescales contributions relative to the explained variance only (i.e., excluding the residual term), and is therefore useful for understanding how the most important terms share the portion of variation that the two-way ORACLE surrogate is actually able to capture.
  3. 3.
    The residual %Total provides a simple diagnostic of how faithfully the two-way decomposition approximates the full ORACLE surrogate: when the residual contribution is small, patterns in the main-effect and interaction plots can be read as describing most of the predictor’s behaviour; when the residual is large, it is safer to focus interpretation on the top few main and interaction terms and to treat the remaining structure as higher-order or model-specific detail.
The main-effect plots and classical interaction plots that follow make these decomposition findings visually explicit by showing how the predicted sound-pressure level changes along each factor and how the most important pairs reinforce or counteract each other across their joint ranges.
Table 9: ANOVA-style decomposition of the ORACLE surrogate on Airfoil (Backbone A).
Source Sum Sq %Total %Expl. Rank Cum%Tot
Frequency 14049.600 18.2 53.4 1 18.2
Angle of attack 2540.306 3.3 9.7 9 21.5
Chord length 3998.431 5.2 15.2 5 26.7
Free stream 1550.353 2.0 5.9 12 28.7
Suction side 8535.002 11.1 32.4 2 39.8
Frequency:Angle of attack 1557.488 2.0 5.9 11 41.8
Frequency:Chord length 5410.700 7.0 20.6 3 48.8
Frequency:Free stream 599.716 0.8 2.3 15 49.6
Frequency:Suction side 4667.970 6.1 17.7 4 55.6
Angle of attack:Chord length 3645.811 4.7 13.9 7 60.4
Angle of attack:Free stream 1398.027 1.8 5.3 13 62.2
Angle of attack:Suction side 3984.059 5.2 15.1 6 67.3
Chord length:Free stream 654.140 0.8 2.5 14 68.2
Chord length:Suction side 2563.853 3.3 9.7 8 71.5
Free stream:Suction side 1691.144 2.2 6.4 10 73.7
Residual 20295.598 26.3 100.0
Total 46611.280 100.0
@@
evaluated on the bin centers used by ORACLE. Several patterns are immediately visible. First, Frequency and Suction side exhibit the largest vertical ranges, matching their dominant sums of squares in Table 9: increasing the frequency from low to high values steadily reduces the predicted sound-pressure level, while larger suction-side displacement thickness is associated with a pronounced decrease in SPL after a sharp peak at very small values. Second, Chord length also shows a mostly monotone decreasing trend, whereas Free stream shows a mild increasing trend over its range. Finally, Angle of attack displays a weakly curved, non-monotone pattern with a broad maximum around moderate angles, suggesting a smaller but nontrivial main effect.
From a usage perspective, we find the following simple heuristics helpful when reading Figure 2:
@@
    Compare the vertical range in each panel; larger ranges indicate factors with stronger main effects and should roughly agree with the ordering in the ANOVA table. This provides a quick visual sanity check that the surrogate decomposition is consistent.
@@
    Inspect the shape of each curve: near-linear trends suggest approximately additive, monotone effects that are easy to reason about and to intervene on (e.g., “lower frequency $\Rightarrow$ lower SPL”), whereas pronounced curvature or non-monotonicity (as for Angle of attack and Suction side) reveals regimes where the effect changes sign or saturates.
@@
    Use the bin centers on the $x$-axis as candidate design points for one-factor-at-a-time interventions: by moving along the curve in Figure 2, practitioners can anticipate how much reduction in SPL is attainable by adjusting a single factor while keeping the others near their empirical distribution.
These main-effect plots thus complement the ANOVA-style table: the table ranks factors by overall contribution, while Figure 2 makes the direction, strength, and nonlinearity of each effect visually explicit.
@@
While Figure 2 visualises how each factor acts on its own, Figure 3 shows how pairs of factors jointly shape the ORACLE surrogate on Airfoil. Each panel in the upper-triangular grid corresponds to a pair $(X_{j},X_{k})$, with the horizontal axis indexing bins of the column factor $X_{k}$ and the three lines representing low, medium, and high bins (Bin 0, Bin 2, Bin 4) of the row factor $X_{j}$. Non-additive structure appears as non-parallel or crossing lines: when the lines remain roughly parallel, changing the row factor shifts the response up or down by an almost constant offset, whereas clearly diverging slopes indicate that the effect of $X_{k}$ depends strongly on the level of $X_{j}$.
Consistent with the ANOVA table in Table 9, several panels in Figure 3 display pronounced interactions. For example, the Frequency:Chord length and Frequency:Suction side panels show that the SPL response to Chord length or Suction side can either increase or decrease depending on whether frequency is in a low or high bin, with the three lines clearly spreading out and sometimes reversing order across the $x$-axis. In contrast, weaker interaction pairs such as those involving Free stream often exhibit nearly parallel lines: the response level changes with the row factor, but the qualitative shape along the column factor is preserved, which is closer to an additive effect.
@@
    Start from the pairs with large interaction sums of squares in Table 9 and verify that their panels show clearly non-parallel or crossing lines. This cross-check ensures that the numerical interaction ranking aligns with visually interpretable patterns.
@@
    Within a given panel, treat the three lines (Bin 0, 2, 4) as a coarse “design grid”: steep changes in slope or reversals between lines mark regimes where simultaneous adjustment of both factors can yield larger gains than tuning either factor alone.
@@
Together, Figure 2 and Figure 3 provide a DoE-style summary of the ORACLE surrogate: the former highlights which single factors matter most and in which direction, while the latter localises where in the input space pairs of factors reinforce or counteract each other.
@@ Appendix E Latent Extensions with ResNet and BERT
Tables 10 and 11 summarize interaction metrics on latent tabular features extracted by ResNet-18 from CIFAR-10. On this image-derived representation, SHAP-family interaction methods consistently achieve higher NDCG@5 and Xfer-NDCG@5 than ORACLE, and also obtain larger CCC and IG@K=5,B=3, indicating better alignment with the target ANOVA interaction strengths and greater best-case intervention gains relative to the ANOVA-based target. The confidence intervals in Table 11 show that these gaps are substantial and robust, with several SHAP variants reaching the upper bound of IG@K=5,B=3 induced by the target interaction maps. In contrast, ORACLE remains comparatively competitive in hotspot overlap (Peak-IoU@0.10), suggesting that its spatial localization of interaction hotspots is less degraded than its ranking and scale estimates in this image-latent regime.
@@
Tables 12 and 13 report interaction metrics on latent tabular features extracted from BERT-base embeddings of AG News. Similar to the ResNet-18 / CIFAR-10 case, SHAP-family interaction methods generally attain higher NDCG@5 and Xfer-NDCG@5 than ORACLE, and several baselines (in particular KS-Int) achieve markedly larger CCC values, indicating better alignment with the scale of the target ANOVA interaction strengths. ORACLE again remains most competitive in hotspot localization: it attains the highest Peak-IoU@0.10 among all methods, suggesting that its spatial interaction maps still track the hotspots of the target ANOVA maps relatively well even when its ranking and scale estimates are weaker. The confidence intervals in Table 13 show that these patterns are robust, with most SHAP variants overlapping near-perfect detection while ORACLE maintains consistently strong Peak-IoU@0.10.

### v3

@@ A.5 Interaction strength consistency and Top-$K$ / NDCG@K alignment
We finally connect Top-$K$ recovery to the ranking metric NDCG@K. Let $G^{\star}_{jk}$ denote the oracle ”gain” associated with pair $(j,k)$, for example $G^{\star}_{jk}=S^{\star}_{jk}$ or a monotone transform thereof. Given a permutation $\pi$ of the pairs, the (unnormalized) DCG@K of $\pi$ with respect to $G^{\star}$ is
@@ A.7 Further theoretical remarks: grid resolution, misspecification, and higher-order terms
All oracle quantities in Appendix A.2 are defined as orthogonal projections in the Hilbert space $L^{2}(P_{X})$ of the trained predictor $f$. In particular, the main-effect and interaction components $(m_{j},g_{jk})$ and the residual $r$ are defined by their $L^{2}(P_{X})$-best approximation properties, regardless of whether $f$ arises from a ”correctly specified” probabilistic model.
@@
  • When the true mechanism contains higher-order or non-factorizable structure, these contributions are distributed across the pairwise surfaces $g_{jk}$ and the residual $r$ in the way that minimizes $L^{2}(P_{X})$ error. Our oracle is therefore ”model-centric”: it asks how $f$ can be best approximated by main and pairwise terms under the observed $P_{X}$, rather than asserting that $g_{jk}$ coincides with a causal or structural notion of two-way interaction.
@@ Appendix B Additional implementation details
After fitting the factorial surrogate in an ANOVA-like basis, we apply a lightweight post-processing step to enforce standard identifiability (sum-to-zero) constraints on the discrete effect tables (main effects and pairwise interactions). Concretely, we (i) center each main-effect table to have zero mean over its levels, (ii) recenter each pairwise interaction table so that all row and column sums vanish, and (iii) absorb any constant offsets induced by these operations into the global mean $\hat{\mu}$ ($\mu$-rebalancing”). The procedure is summarized in Algorithm 1; it requires only a few passes over the discrete tables and is negligible compared to training the backbone or fitting the surrogate.
Algorithm 1 Recentering and $\mu$-rebalancing for discrete effect tables (identifiability normalization)
@@
1:// Center main effects (zero mean over levels)
@@
11:  $\bar{g}_{jk}\leftarrow\frac{1}{|\mathcal{B}_{j}||\mathcal{B}_{k}|}\sum_{\ell_{j}\in\mathcal{B}_{j}}\sum_{\ell_{k}\in\mathcal{B}_{k}}\hat{g}_{jk}(\ell_{j},\ell_{k})$
@@
14:  subtract row means so that $\sum_{\ell_{j}\in\mathcal{B}_{j}}\hat{g}_{jk}(\ell_{j},\ell_{k})=0$ for all $\ell_{k}\in\mathcal{B}_{k}$
15:  subtract column means so that $\sum_{\ell_{k}\in\mathcal{B}_{k}}\hat{g}_{jk}(\ell_{j},\ell_{k})=0$ for all $\ell_{j}\in\mathcal{B}_{j}$
@@
17: Normalized $\hat{\mu}$, $\{\hat{m}_{j}\}$, $\{\hat{g}_{jk}\}$ satisfying identifiability (sum-to-zero) constraints
@@ Appendix D Classical Main-Effect and Interaction Plots
evaluated at the bin centers used by ORACLE. Several patterns are immediately visible. First, Frequency and Suction side exhibit the largest vertical ranges: increasing frequency from low to high values steadily reduces the predicted sound-pressure level, while larger suction-side displacement thickness is associated with a pronounced decrease in SPL after a sharp peak at very small values. Second, Chord length shows a largely monotone decreasing trend, whereas Free stream shows a mild increasing trend over its range. Finally, Angle of attack displays a weakly curved, non-monotone pattern with a broad maximum at moderate angles, suggesting a smaller but nontrivial main effect.
From a practical perspective, we find the following simple heuristics helpful when reading Figure 2:
@@
    Compare the vertical range across panels; larger ranges indicate stronger main effects and provide a quick visual prioritization.
@@
    Inspect the shape of each curve: near-linear trends suggest approximately additive, monotone effects that are easy to reason about and to intervene on (e.g., ”lower frequency $\Rightarrow$ lower SPL”), whereas pronounced curvature or non-monotonicity (as for Angle of attack and Suction side) reveals regimes where the effect changes sign, peaks, or saturates.
@@
    Use the bin centers on the $x$-axis as candidate design points for one-factor-at-a-time interventions: by moving along the curve, practitioners can anticipate how much reduction in SPL is attainable by adjusting a single factor while keeping the others near their empirical distribution.
Overall, the main-effect plots provide a compact summary of the direction, magnitude, and nonlinearity of each marginal effect, and help identify promising targets for simple interventions.
@@
While Figure 2 summarizes how each factor acts on its own, Figure 3 shows how pairs of factors jointly shape the ORACLE surrogate on Airfoil. Each panel in the upper-triangular grid corresponds to a pair $(X_{j},X_{k})$, with the horizontal axis indexing bins of the column factor $X_{k}$ and the three lines representing low, medium, and high bins (Bin 0, Bin 2, Bin 4) of the row factor $X_{j}$. Non-additive structure appears as non-parallel or crossing lines: when the lines remain roughly parallel, changing the row factor shifts the response by an almost constant offset, whereas clearly diverging slopes indicate that the effect of $X_{k}$ depends strongly on the level of $X_{j}$.
Several panels in Figure 3 display pronounced interactions. For example, the Frequency:Chord length and Frequency:Suction side panels show that the SPL response to Chord length or Suction side can strengthen, weaken, or even reverse depending on whether Frequency is in a low or high bin, with the three lines clearly spreading out and sometimes changing order across the $x$-axis. In contrast, weaker pairs (often those involving Free stream) tend to exhibit nearly parallel lines: the response level shifts with the row factor, but the qualitative shape along the column factor is largely preserved, which is closer to an additive effect.
@@
    Look first for panels with clearly non-parallel or crossing lines; these identify pairs where joint adjustment can matter more than one-factor-at-a-time tuning.
@@
    Within a given panel, treat the three lines (Bin 0, Bin 2, Bin 4) as a coarse ”design grid”: steep slope changes and line crossings mark regimes where the effect of one factor depends strongly on the level of the other.
@@
Together, Figure 2 and Figure 3 provide a DoE-style summary of the ORACLE surrogate: the former highlights which single factors matter most and in which direction, while the latter localizes where in the input space pairs of factors reinforce or counteract each other.
@@ Appendix E Latent Extensions with ResNet and BERT
Tables 9 and 10 summarize interaction metrics on latent tabular features extracted by ResNet-18 from CIFAR-10. On this image-derived representation, SHAP-family interaction methods consistently achieve higher NDCG@5 and Xfer-NDCG@5 than ORACLE, and also obtain larger CCC and IG@K=5,B=3, indicating better alignment with the target ANOVA interaction strengths and greater best-case intervention gains relative to the ANOVA-based target. The confidence intervals in Table 10 show that these gaps are substantial and robust, with several SHAP variants reaching the upper bound of IG@K=5,B=3 induced by the target interaction maps. In contrast, ORACLE remains comparatively competitive in hotspot overlap (Peak-IoU@0.10), suggesting that its spatial localization of interaction hotspots is less degraded than its ranking and scale estimates in this image-latent regime.
@@
Tables 11 and 12 report interaction metrics on latent tabular features extracted from BERT-base embeddings of AG News. Similar to the ResNet-18 / CIFAR-10 case, SHAP-family interaction methods generally attain higher NDCG@5 and Xfer-NDCG@5 than ORACLE, and several baselines (in particular KS-Int) achieve markedly larger CCC values, indicating better alignment with the scale of the target ANOVA interaction strengths. ORACLE again remains most competitive in hotspot localization: it attains the highest Peak-IoU@0.10 among all methods, suggesting that its spatial interaction maps still track the hotspots of the target ANOVA maps relatively well even when its ranking and scale estimates are weaker. The confidence intervals in Table 12 show that these patterns are robust, with most SHAP variants overlapping near-perfect detection while ORACLE maintains consistently strong Peak-IoU@0.10.